OpenStudy (anonymous):

The product of three prime numbers is seven times their sum. Find the three prime numbers.

OpenStudy (anonymous):

If we let the numbers as: x, y and z Then : xyz = 7(x + y + z) And @hartnn is right.. Ha haha..

hartnn (hartnn):

shortest way is to do trials, i started with 2,3,5 then 3,5,7 and i got it :P

OpenStudy (experimentx):

xyz = 7(x + y + z) xyz - 7z = 7(y+x) z(xy -7) = 7(y+x) ----------------- i don't have a complete logic here z=7, problem reduces to xy -7 = y+x

OpenStudy (anonymous):

And after that you will use Hit and Trial Method again??

OpenStudy (experimentx):

only physicist are allowed to do that.

OpenStudy (anonymous):

So, what will you use if you are not a physicist ??

OpenStudy (experimentx):

I'm half physicist .. :P

OpenStudy (anonymous):

Then I think @hartnn is full physicist.. Ha ha ha..

OpenStudy (experimentx):

xy = y+x +7 y(x-1) = x+7 y = (x+7)/(x-1) <--- this should be an integer (prime ) maybe we could impose this constrain here.

OpenStudy (experimentx):

mukushla is perfect for this kind of problem ... where did he go?

OpenStudy (anonymous):

Yes I am looking for him too..

OpenStudy (anonymous):

He is offline but we should mention his name @mukushla

OpenStudy (experimentx):

it is easy to see that 2,3 will satisfy this criteria (x+7)/(x-1) = 2+7 = 9 <-- not a prime (x+7)/(x-1) = 10/2 = 5

OpenStudy (anonymous):

There if x is 3 then we get y = 5 And they both are prime and anyhow according to your logic z = 7 too.

OpenStudy (experimentx):

this isn't the problem ... you have to prove either this is infinite solutions (of some form) /// some finite solution /// or just unique solution.

OpenStudy (anonymous):

Yeah may be we have other prime numbers which will satisfy the given condition.. Or may not be..

hartnn (hartnn):

maybe the word 'consecutive' prime numbers, in the question can help simplify the solution....

OpenStudy (experimentx):

also there is hidden logic here (probably not nice) z(xy -7) = 7(y+x) ----------------- i don't have a complete logic here z=7,

OpenStudy (anonymous):

But there isn't any word named "Consecutive" there..

OpenStudy (anonymous):

The hidden logic: xyz = 7 (x+y+z) Since a, b and c are prime, so, one of a, b, c should be 7

OpenStudy (anonymous):

Oh wow hidden but not complete logic..

OpenStudy (anonymous):

x, y and z will be multiple of 7 if one out of them is multiple of 7 or 7 itself..

OpenStudy (anonymous):

It can't be multiple of 7 since it is a PRIME!

OpenStudy (anonymous):

of course except 7 :P

OpenStudy (anonymous):

And x and y and z are prime too.. I am saying according to the equation @RolyPoly

OpenStudy (anonymous):

According to this equation: xyz = 7(x + y + z) Product of x, y and z will be multiple of 7 so, one out of x, y and z should be 7..

OpenStudy (anonymous):

Put x = 7, then yz = 7 +y +z So, z = (y+7) / (y-1) We must use trial and error here? Or is there any other better way?

OpenStudy (anonymous):

If we use the word consecutive here, then what are the prime numbers consecutive to each other?? 2 3 and 5 or 3 5 and 7 Not more than that I think @hartnn

OpenStudy (anonymous):

Any othere better way @mukushla can tell you if any.. But for now, Hit and Trial method is what you can do..

OpenStudy (anonymous):

so one of these must be seven because xyz is divisible by 7 so say x=7 \[yz=7+y+z\]\[y=\frac{z+7}{z-1}=\frac{z-1+8}{z-1}=1+\frac{8}{z-1}\]so 8 is divisible by z-1 so\[z-1=1,2,4,8\]it gives\[z=2,3,5\]we are done

OpenStudy (anonymous):

Great..